### v2

@@
Abstract
Deduplication is a major focus for assembling and curating training datasets for large language models (LLM) – detecting and eliminating additional instances of the same content – in large collections of technical documents. Unrestrained, duplicates in the training dataset increase training costs and lead to undesirable properties such as memorization in trained models or cheating on evaluation. Contemporary approaches to document-level deduplication are often extremely expensive in both runtime and memory. We propose LSHBloom, an extension to MinhashLSH, which replaces the expensive LSHIndex with lightweight Bloom filters. LSHBloom demonstrates the same deduplication performance as MinhashLSH with only a marginal increase in false positives (as low as 1e-5 in our experiments); demonstrates competitive runtime (270% faster than MinhashLSH on peS2o); and, crucially, uses just 0.6% of the disk space required by MinhashLSH to deduplicate peS2o. We demonstrate that this space advantage scales with increased dataset size—at the extreme scale of several billion documents, LSHBloom promises a 250% speedup and a 54$\times$ space advantage over traditional MinHashLSH scaling deduplication of text datasets to many billions of documents.
marginparsep has been altered.
topmargin has been altered.
marginparwidth has been altered.
marginparpush has been altered.
The page layout violates the ICML style. Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
LSHBloom: Memory-efficient, Extreme-scale Document Deduplication
Anonymous Authors1
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Scholarly publications and related literature represent a tremendous body of often carefully encoded human knowledge. Google Scholar is estimated to index some 400 million documents Gusenbauer (2019), suggesting that the total number of unique documents worldwide numbers at least in the hundreds of millions. And a growing number of digital repositories mean that researchers can today access millions of such documents with ease, and with effort, tens of millions or more.
@@ 1 Introduction
Developments in large language models (LLMs) suggest that such corpora can be used to develop powerful agents able to assist with technical work in a variety of fields Zhang et al. (2024). However, seizing this opportunity at the largest scale requires addressing challenging deduplication problems. Any large corpus of technical documents assembled from diverse locations will include multiple copies of many documents: not only identical copies but also copies that vary, subtly or otherwise, for reasons such as different publication formats, the use of different parsers, copyediting, and versioning. As is well known, duplication in training datasets can lead to a host of undesirable behaviors, from excessive training times to predictive biases and training data memorization Lee et al. (2021) in trained models. Text duplication between training and evaluation sets also hinders thorough, unbiased assessments of model performance, with much evidence suggesting that standard open-source benchmarks have leaked into LLM training sets Chandran et al. (2024), illegitimately increasing perceived model performance by 8–11% Dekoninck et al. (2024).
@@
These considerations motivate the problem that we tackle in this paper, namely the accurate and efficient deduplication of large numbers (10s to 100s of millions) of scholarly and technical documents. We approach this problem from two perspectives. First, we demonstrate via careful evaluation that existing state-of-the-art approaches to the deduplication of technical documents suffer from poor deduplication performance (i.e., low precision and/or recall) and/or excessive computational costs. Due to the frequent loss of metadata and variation in documents, deduplication of technical documents requires approximate textual comparison methods such as Jaccard similarity, which may be accelerated via the approximating MinHashing method Broder et al. (1998). The infeasibility of pairwise comparisons in large datasets motivated the development of the MinHashLSH method Leskovec et al. (2014), which applies a locality-sensitive hashing (LSH) scheme to MinHash signatures to avoid comparing dissimilar documents. However, the associated MinHashLSH index can grow extremely large. For example, applying MinHashLSH to the peS2o dataset Soldaini & Lo (2023) of just 8M full-text and 31M abstracts takes more than 14 hours on a single 32-core node of the Polaris supercomputer and requires more than 300GB of disk space (see Section 5.3.1).
Second, we introduce a new method, LSHBloom, that improves significantly on the state of the art in terms of the computational costs required to achieve a desired level of deduplication performance. The key insight here is that the process that MinHashLSH employs to find matches between MinHash signatures can be approximated, with a modest and controllable loss of accuracy, by using Bloom filters. This use of Bloom filters enables LSHBloom to run almost three times faster than MinHashLSH by interfacing with inexpensive bit arrays rather than inserting and querying against the comparatively slow tree data structures that comprise the traditional LSHIndex. Most importantly, this use of Bloom filters means that LSHBloom requires just 0.6% of the disk space consumed by the traditional MinHashLSH index to deduplicate peS2o Soldaini & Lo (2023), a dataset of 39 million documents. We demonstrate that this space advantage remains as the number of documents increases and makes it feasible for LSHBloom to process several billion documents with space savings of 54$\times$ relative to MinHashLSH without sacrificing deduplication quality.
2 Background
@@
2.1 Defining the Deduplication Problem
Text deduplication algorithms seek to identify and remove duplicate instances of the same or similar content from a document, dataset, or text corpus. These methods can be applied at different resolutions (e.g., to documents, paragraphs, sentences, or even tokens) and scopes (e.g., across a corpus or within a single document). Deduplication methods may also differ in how they define duplication (e.g., exact matches, n-gram overlap, “fuzzy”). In evaluating a deduplication method for a particular application, we are concerned with its precision, recall, and computational performance. Low precision, i.e., excessive false positives, is harmful as it unnecessarily reduces the size of the training dataset and limits an LLM’s ability to learn the underlying distribution of the text data. Low recall, i.e., excessive false negatives, admits a great degree of duplicate content in our dataset, with negative effects on learning as well as potentially inducing degenerate behavior such as memorization. High computational performance is critical due to the extreme size of modern LLM datasets. For example, peS2o Soldaini & Lo (2023) contains roughly 39 million documents, Google Scholar is estimated to index some 400 million academic papers and other scholarly literature Gusenbauer (2019), DOLMA Soldaini et al. (2024) contains roughly 5B documents, and RedPajama Computer (2023) has just over 100B documents—scales at which naive pairwise comparisons become infeasible. Text deduplication algorithms are needed that can efficiently process billions of documents while maintaining high precision and recall.
@@ 2.1 Defining the Deduplication Problem
Just how to define duplication is a subtle question. For example, one could stipulate simply that two documents must be byte-for-byte identical: equivalent to setting $S$ to measure byte overlap and $T$ = 1.0 above. However, this definition may exclude many document pairs that we would like to consider duplicates due to great similarity in their token distributions or semantic content, as are encountered frequently in practical text ingestion pipelines. For example, the same document may be available from two separate venues (say, a pre-print on Arxiv and a camera-ready conference publication) that differ only slightly due to editorial decisions. When web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and various format-specific text parsing pipelines tend to induce small differences even when applied to the same underlying document. Even simply parsing two PDFs of the same document introduces complications—for instance, the choice of PDF parser may be influenced by the document’s origin, publisher, or language. PDF parsers that rely on optical character recognition (OCR) can also introduce artifacts that render exact matching woefully ineffective.
@@
For these reasons, many deduplication techniques choose to take a “fuzzy” perspective on duplication, allowing for minor differences between items in duplicate pairs. One way to do this is to alter one’s choice of similarity metric to encapsulate a broader semantic notion of similarity. This is common in so-called semantic deduplication techniques such as SemDeDup Abbas et al. (2023), which employ embeddings from pre-trained language models to identify semantically similar text data. Unfortunately, these techniques require training language models and performing inference on each data point before clustering similar items to identify duplicates, which is prohibitively expensive in the extreme-scale regime with billions of documents. It is more efficient to instead establish a similarity metric based on the raw text content of each document.
2.2 Jaccard Similarity and MinHash Approximation
@@
Popular techniques such as those used to build the DOLMA Soldaini et al. (2024) and DataComp-LM Li et al. (2024) datasets measure n-gram overlap between documents to identify duplicates. Alternatively, some techniques—such as MinHashLSH—choose a set-theoretic notion of document similarity.
@@
A text document can be viewed as a bag of n-grams, and the similarity between two such sets, $A$ and $B$, defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
@@
However, computing Jaccard similarity requires evaluating set membership for each item present in the union of both sets. For documents, this set size can be large, and computing the Jaccard similarity explicitly can be time-consuming. The MinHashing procedure Broder et al. (1998) provides a means of efficiently approximating the Jaccard similarity of two documents. Rather than compare the elements of each set directly, this procedure randomly permutes the order of set elements to produce so-called MinHash signatures. The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets Broder et al. (1998); Leskovec et al. (2014). Therefore, we can repeatedly apply random permutations to the two sets’ elements and compare them for equality to obtain an estimate of the Jaccard similarity. Our estimate grows more accurate as we increase our sample size.
2.3 MinHashLSH for Scalable MinHash
@@
The use of MinHash-estimated Jaccard similarity, $J_{H}$, as an approximation for Jaccard similarity, $J$, provides for more efficient similarity computations because the procedure employed computed to estimate the similarity of two documents has linear complexity in the number of random permutations applied, rather than quadratic complexity in the number of n-grams present in each document. However, the need to compare every pair of documents for similarity during deduplication still introduces a quadratic cost when comparing all documents in a corpus.
@@
Fortunately, there exists a technique that allows us to compare only those pairs with a high likelihood of similarity, rather than all pairs. MinHashLSH Leskovec et al. (2014) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents. Given a set of MinHash signatures, one can construct a MinHash signature matrix in which each column contains all MinHash values for a single document. Then, given a desired Jaccard similarity threshold, $T$, one can determine an optimal band size, $b$, with which to group the rows of the MinHash signature matrix. This results in $b$ groupings of rows (bands) of size $r$. With high probability, if two bands of the MinHash signature matrix are exactly identical for two documents (columns), then those two documents have Jaccard similarity greater than or equal to $T$. This enables us to produce a series of candidate pairs of documents that we can take as duplicate pairs or further evaluate for similarity. This method avoids pairwise comparison between every pair of documents in the corpus but introduces associated false positive and false negative probabilities due to the use of locality-sensitive hashing.
@@
One can implement a prefix-tree or suffix array to search efficiently for candidate duplicate documents when evaluating a new document for duplication. However, this requires storing a representation of each document, and while MinHash representations are small compared to raw documents, the MinHashLSH index becomes large for large datasets: e.g., 23 TB for 5B documents. Additionally, inserting and querying these tree-structured indices constitutes a significant portion (>80%) of algorithm runtime, which can become prohibitive at extreme scales (see Figure 1). Here we remedy this space and runtime issue by employing a novel combination of MinHashLSH and Bloom filters.
@@
Figure 1: Breakdown of wall clock time on 5% of peS2o for the conventional MinHashLSH algorithm and our LSHBloom method
2.4 Bloom Filter
@@
The Bloom filter Bloom (1970) is a space-efficient probabilistic data structure used to test whether an element is a member of a set. It can determine with certainty if an element is not in the set, but if it indicates the element is in the set, there is a small probability of error. That is, the Bloom filter may admit false positives but not false negatives.
@@
Bloom filters have been used extensively in distributed computing for a variety of purposes Broder & Mitzenmacher (2004); Chervenak et al. (2008); Tarkoma et al. (2011).
3 Related Work
Having provided background on efficient document similarity measures, we briefly review related work on deduplication methods employed by groups developing LLMs. LLM developers use various methods, many ad hoc, to deduplicate the large quantities of text content used for LLM training regardless of whether the text comes from a traditional source or synthetic data. Many leading LLMs do not discuss the deduplication and curation methods they use, such as Llama-2 Touvron et al. (2023b), Claude, Gemini Gemini Team (2024a; b), and GPT versions 3 and later OpenAI (2024). Others, such as Palm2 Anil et al. (2023) and Gemma Gemma Team (2024); Gemini Team (2024b), provide only limited details. Others, including several leading models, provide extensive details about how data for their model was dededuplicated including Olmo’s Dolma dataset Soldaini et al. (2024) as well as its successor DataComp-LM Li et al. (2024), Llama 1 Touvron et al. (2023a), Llama 3 Dubey et al. (2024), and OPT Zhang et al. (2022). For example, Llama 3 Dubey et al. (2024) was trained on a dataset that was deduplicated by using MinHash over the entire dataset, line-level deduplication using CCNet for lines replicated over six times in each bucket of text based on coarse categorization (e.g., math reasoning, geometry/trigonometry), and URL deduplication for web data. Phi uses an approach described by Gunasekar et al. (2023) and elaborated on in their Phi 2.0 report Javaheripi (2023) that starts from deduplicated code and text, but they do not discuss deduplicating after generation.

### v3

@@
Abstract.
[…]
PVLDB Reference Format:
PVLDB, 19(1): XXX-XXX, 2025.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
1. Introduction
Modern large language model (LLM) training pipelines require the ingestion of many millions of documents to enable internet-scale language pre-training. These documents are obtained through web scraping, deals with academic institutions, and relationships with publishers. The sheer scale of this data ingestion process presents a substantial data management and processing burden when assembling large-scale text databases for continuous LLM pre-training. LLM performance has also been shown to be sensitive to data quality (alignment of data with downstream tasks) (tan202415pintstechnicalreportpretraining) as well as data redundancy–excessive redundancy can cause performance scaling below the predicted rate of established scaling laws (chen-etal-2025-revisiting). Moreover, per-epoch training costs scale in the number of training data points. As such, it is highly desirable to remove outright uninformative data points from internet-scale text datasets–to alleviate the data management burden, to reduce the costs of pre-training, and to enhance LLM performance.
@@ 1. Introduction
In this work, we focus on reducing data redundancy in text. Duplicate text data–that is, instances of the same content or content that only subtly differ from one another due to differences in publication format, copyediting, and versioning–are ubiquitous in internet-scale text datasets and are well known to lead to a host of undesirable outcomes, from excessive training times to the introduction of predictive biases and training data memorization (lee2021deduplicating), in trained models. Text duplication between training and evaluation sets also hinders thorough, unbiased assessments of model performance, with much evidence suggesting that standard open-source benchmarks have leaked into LLM training sets (Chandran2024PrivateBT), illegitimately inflating perceived model performance by 8–11% (Dekoninck2024EvadingDC).
These considerations motivate us to address the accurate and efficient deduplication of large numbers (100s of millions) of documents when preprocessing internet-scale text datasets.
@@
We demonstrate via careful evaluation that existing state-of-the-art approaches to text deduplication suffer from poor deduplication performance (i.e., low precision and/or recall) and/or excessive computational costs. Due to the frequent loss of metadata and subtle variation in documents, deduplication of technical documents requires approximate textual comparison methods such as Jaccard similarity, which may be accelerated via the approximating MinHashing method (broder1998min). The infeasibility of pairwise comparisons in large datasets motivated the development of MinHashLSH (leskovec2014finding), which applies a locality-sensitive hashing (LSH) scheme to MinHash signatures to avoid comparing dissimilar documents. MinHashLSH is by far the most popular method for text deduplication in LLM data pipelines; however, the associated MinHashLSH index can grow extremely large. For example, applying MinHashLSH to the peS2o dataset (peS2o) of just 39M documents takes more than 35 hours on a single 32-core node of the Polaris supercomputer and requires more than 200 GB of disk space (see subsubsection 5.4.1).
[…]
MinHashLSH is the most popular and performant state-of-the-art text deduplication technique available today; LSHBloom can be employed as a drop-in replacement for MinHashLSH–preserving this superior performance while enabling significant gains in scaling capacity as we increase our training dataset sizes. We know from historical data that LLM training dataset sizes will almost quadruple each year (epoch2024datasetsizetrend). LSHBloom presents an opportunity to create robust data processing infrastructure for LLM pre-training that is effective, efficient, and scalable well into the future. In summary, the contributions of our work are the following:
  1. (1)
    We present the first side-by-side benchmark comparison of state-of-the-art text deduplication algorithms utilized LLM data processing pipelines, comparing both their deduplication performance (in terms of precision, recall, and F1) as well as their resource requirements as we increase the size of the underlying text data
  2. (2)
    We present the first study on the effects of tuning the hyperparameters for each of these algorithms, and present the best parameter settings for each
  3. (3)
    We develop LSHBloom, a novel text deduplication algorithm that outperforms or matches the existing state-of-the-art in deduplication performance while requiring significantly fewer resources and therefore enhancing scalability for the internet-scale data processing regime
2. Background
@@
2.1. Defining the Deduplication Problem
Text deduplication algorithms seek to identify and remove duplicate instances of the same or similar content from a document, dataset, or text corpus. These methods can be applied at different resolutions (e.g., to documents, paragraphs, sentences, or even tokens) and scopes (e.g., across a corpus or within a single document). Deduplication methods may also differ in how they define duplication (e.g., exact matches, n-gram overlap, “fuzzy”). In evaluating a deduplication method for a particular application, we are concerned with its precision, recall, and computational performance. Low precision, i.e., excessive false positives, is harmful as it unnecessarily reduces the size of the training dataset and limits an LLM’s ability to learn the underlying distribution of the text data. Low recall, i.e., excessive false negatives, admits a great degree of duplicate content in our dataset, with negative effects on learning as well as potentially inducing degenerate behavior such as memorization. High computational performance is critical due to the extreme size of modern LLM datasets. For example, peS2o (peS2o) contains roughly 39 million documents, Google Scholar is estimated to index some 400 million academic papers and other scholarly literature (gusenbauer2019google), DOLMA (soldaini_dolma_2024) contains roughly 5B documents, and RedPajama (together2023redpajama) has just over 100B documents—scales at which naive pairwise comparisons become infeasible. Text deduplication algorithms are needed that can efficiently process billions of documents while maintaining high precision and recall.
@@ 2.1. Defining the Deduplication Problem
Just how to define duplication is a subtle question. For example, one could stipulate simply that two documents must be byte-for-byte identical: equivalent to setting $S$ to measure byte overlap and $T$ = 1.0 above. However, this definition may exclude many document pairs that we would like to consider duplicates due to great similarity in their token distributions or semantic content, as are encountered frequently in practical text ingestion pipelines. For example, the same document may be available from two separate venues (say, a pre-print on Arxiv and a camera-ready conference publication) that differ only slightly due to editorial decisions. When web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and the appropriate format-specific text parsing pipelines will induce small differences even when applied to the same underlying document. Even simply parsing two PDFs of the same document introduces complications—for instance, the choice of PDF parser may be influenced by the document’s origin, publisher, or language. PDF parsers that rely on optical character recognition (OCR) can also introduce artifacts that render exact matching woefully ineffective.
@@
For these reasons, many contemporary deduplication techniques take a ”fuzzy” perspective on duplication, allowing for minor differences between items in duplicate pairs. One way to do this is to alter one’s choice of similarity metric to encapsulate a broader notion of similarity. This is common in so-called semantic deduplication techniques such as SemDeDup (abbas2023semdedup), which employ embeddings from pre-trained language models to identify semantically similar text data. Unfortunately, these techniques require training language models and performing inference on each data point before clustering similar items to identify duplicates, which is prohibitively expensive in the extreme-scale regime with billions of documents. It is more efficient to instead establish a similarity metric based on the raw text content of each document.
2.2. Jaccard Similarity and MinHash Approximation
@@
Popular techniques such as those used to build the DOLMA (soldaini_dolma_2024) and DataComp-LM (li2024datacomp) datasets measure n-gram overlap between documents to identify duplicates. Alternatively, some techniques—such as MinHashLSH—choose a set-theoretic notion of document similarity.
@@
In this paradigm, a text document is viewed as a bag of n-grams, and the similarity between two such sets, $A$ and $B$, defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
@@
However, computing Jaccard similarity requires evaluating set membership for each item present in the union of both sets. For documents, this set size can be large, and computing the Jaccard similarity explicitly can be time-consuming. The MinHashing procedure (broder1998min) provides a means of efficiently approximating the Jaccard similarity of two documents. Rather than compare the elements of each set directly, this procedure randomly permutes the order of set elements to produce so-called MinHash signatures. The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets (broder1998min; leskovec2014finding). Therefore, we can repeatedly apply the MinHashing procedure to any two documents and compare the resulting hashes for equality to obtain an estimate of their Jaccard similarity. This estimate grows more accurate as we increase our sample size.
2.3. MinHashLSH for Scalable MinHash
@@
The use of MinHash-estimated Jaccard similarity, $J_{H}$, as an approximation for Jaccard similarity, $J$, provides for more efficient similarity computations because the procedure employed to estimate the similarity of two documents has linear complexity in the number of random permutations applied, rather than quadratic complexity in the number of n-grams present in each document. However, the need to compare every pair of documents for similarity during deduplication still introduces quadratic cost-scaling with respect to the number of documents in our dataset.
@@
Fortunately, there exists a technique that allows us to compare only those pairs with a high likelihood of similarity, rather than all pairs. MinHashLSH (leskovec2014finding) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents. Given a set of MinHash signatures, one can construct a MinHash signature matrix in which each column contains all MinHash values for a single document. Then, given a desired Jaccard similarity threshold, $T$, one can determine an optimal band size, $b$, with which to group the rows of the MinHash signature matrix. This results in $b$ groupings of rows (bands) of size $r$. With high probability, if two bands of the MinHash signature matrix are exactly identical for two documents (columns), then those two documents have Jaccard similarity greater than or equal to $T$. This procedure enables us to produce a series of candidate pairs of documents that we can take as duplicate pairs or further evaluate for similarity. This method avoids pairwise comparison between every pair of documents in the corpus but introduces associated false positive and false negative probabilities due to the use of locality-sensitive hashing.
@@
One can implement a prefix-tree or suffix array to search efficiently for candidate duplicate documents when evaluating a new document for duplication. However, this requires storing a representation of each document, and while MinHash representations are small compared to raw documents, the MinHashLSH index becomes large for large datasets: e.g., 277 TB for 5B documents. Additionally, inserting and querying these tree-structured indices constitutes a significant portion (¿85%) of algorithm runtime, which can become prohibitive at extreme scales (see Figure 1). Here we remedy this space and runtime issue by employing a novel combination of MinHashLSH and Bloom filters.
@@
Figure 1. Breakdown of wall clock time on 10% of peS2o for the conventional MinHashLSH algorithm and our LSHBloom method
2.4. Bloom Filter
@@
The Bloom filter (bloom1970space) is a space-efficient probabilistic data structure used to test whether an element is a member of a set. It can determine with certainty if an element is not in the set, but if it indicates the element is in the set, there is a small probability of error. That is, the Bloom filter may admit false positives but not false negatives.
@@
Bloom filters have been used extensively in distributed computing for a variety of purposes (broder2004network; chervenak2008globus; tarkoma2011theory).
3. Related Work
Having provided background on efficient document similarity measures, we briefly review related work on deduplication methods employed by groups developing LLMs. LLM developers use various methods, many ad hoc, to deduplicate the large quantities of text content used for LLM training regardless of whether the text comes from a traditional source or synthetic data. Many leading LLMs do not discuss the deduplication and curation methods they use, such as Llama-2 (touvron_llama_2023-1), Claude, Gemini (gemini_team_gemini_2024; gemini_team_gemini_2024-1), and GPT versions 3 and later (openai_gpt-4_2024). Others, such as Palm2 (anil2023palm) and Gemma (gemma_team_gemma_2024; gemini_team_gemini_2024-1), provide only limited details. Others, including several leading models, provide extensive details about how data for their model was deduplicated including Olmo’s Dolma dataset (soldaini_dolma_2024) as well as its successor DataComp-LM (li2024datacomp), Llama 1 (touvron_llama_2023), Llama 3 (dubey_llama_2024), and OPT (zhang_opt_2022). For example, Llama 3 (dubey_llama_2024) was trained on a dataset that was deduplicated by using MinHash over the entire dataset, line-level deduplication using CCNet for lines replicated over six times in each bucket of text based on coarse categorization (e.g., math reasoning, geometry/trigonometry), and URL deduplication for web data. Phi uses an approach described by gunasekar_textbooks_2023 and elaborated on in their Phi 2.0 report (javaheripi_surprising_2023) that starts from deduplicated code and text, but they do not discuss deduplicating after generation. C4 (t5_c4_raffel2020) applies only heuristic filtering for data quality (e.g., removing lines without punctuation or fewer than three words). RedPajama (together2023redpajama) uses both exact matching via Bloom filters and fuzzy deduplication via MinHashLSH. Datasets such as the Pile (gao2020pile800gbdatasetdiverse) and more recent releases such as FineWeb2 (penedo2025fineweb2pipelinescale) apply MinHashLSH for fuzzy deduplication.